OpenStudy (anonymous):

A = P(1+i)^n 1. Isolate the variable P 2. Isolate the variable i

OpenStudy (anonymous):

where did you get 10 from?

OpenStudy (anonymous):

this is so confusing :(

OpenStudy (anonymous):

Polar form...?

myininaya (myininaya):

take natural log of both sides \[\ln(A)=\ln(P(1+i)^n)=\ln(P)+\ln(1+i)^n=\ln(P)+n*\ln(1+i)\] so we have \[\ln(A)=\ln(P)+n*\ln(1+i)\] subract ln(P) on both sides \[\ln(A)-\ln(P)=n*\ln(1+i)\] remember ln(x/y)=lnx-lny so we can write \[\ln(\frac{A}{P})=n*\ln(1+i)\] now multiply both sides by 1/n \[\frac{1}{n}*\ln(\frac{A}{P})=\frac{1}{n}*n*\ln(1+i)\] \[\frac{1}{n}\ln(\frac{A}{P})=\ln(1+i)\] \[\ln(\frac{A}{P})^\frac{1}{n}=\ln(1+i)\] do e^ on both sides \[(\frac{A}{P})^\frac{1}{n}=1+i\] subtract i on both sides \[(\frac{A}{P})^\frac{1}{n}-1=i\] but this way is probably way more confusing (and unnecessary) lol so we have \[A=P(1+i)^n\] we want to solve for i so first divide P on both sides \[\frac{A}{P}=(1+i)^n\] now do ^(1/n) on both sides \[(\frac{A}{P})^\frac{1}{n}=(1+i)^{n*(\frac{1}{n})}\] \[(\frac{A}{P})^\frac{1}{n}=(1+i)^1\] \[(\frac{A}{P})^\frac{1}{n}=1+i\] now subtract 1 on bboth sides \[(\frac{A}{P})^\frac{1}{n}-1=i\]

myininaya (myininaya):

you probably don't really want to look at the first way i did but the second way she be pretty easy to understand what do you think?

OpenStudy (anonymous):

i think the way you did it is confusing, because here on the question theyre just asking to isolate the variable for P so wouldnt i do something along the lines of P=(a)(1+i)^n

myininaya (myininaya):

the second way?

OpenStudy (anonymous):

yes

myininaya (myininaya):

you want to solve for i we need to work on getting i by itself so we notice i is inside here: (1+i)^n right? but (1+i)^n is being multiplied by that P so to get to (1+i)^n by itself we need to divide both sides by P first

myininaya (myininaya):

\[\frac{A}{P}=(1+i)^n\] so you understand this part right?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok but i is inside that paranthese that is to the n power to get rid of the n power we need to raise both sides by 1/n becasue we know n*1/n will give us 1 right? \[(\frac{A}{P})^\frac{1}{n}=((1+i)^{n})^\frac{1}{n}\] do you understand this part? do you undestand why I raised both sides to the 1/n power?

OpenStudy (anonymous):

remember im trying to isolate the variable by themselves so im trying to isolate i by itself and p by its self using the formula A = P( 1+ i )^n

myininaya (myininaya):

since n*1/n=1 we have \[(\frac{A}{P})^\frac{1}{n}=(1+i)^1\]

myininaya (myininaya):

but \[(1+i)^1=(1+i)=1+i\]

myininaya (myininaya):

so we have \[(\frac{A}{P})^\frac{1}{n}=1+i\] to isolate i by itself now we need to subtract 1 on both sides

myininaya (myininaya):

\[(\frac{A}{P})^\frac{1}{n}-1=i\]

myininaya (myininaya):

if you don't like the way this is written you can always write it going the other way: \[i=(\frac{A}{P})^\frac{1}{n}-1\]

myininaya (myininaya):

they mean the samethings

OpenStudy (anonymous):

It also works out if u do 1+i as sqrt(2) and pi/4 plus complex powers.

OpenStudy (anonymous):

But your way is simpler...(much)

myininaya (myininaya):

lol wedphill do you understand now?

OpenStudy (anonymous):

yes i think i get it now

OpenStudy (anonymous):

thank you so much for your help :)

myininaya (myininaya):

you aren't just saying that to get rid of me right?

OpenStudy (anonymous):

nooo

myininaya (myininaya):

lol ok im glad i can help

OpenStudy (anonymous):

I was just confused at the part when i had to isolate i

OpenStudy (anonymous):

i didnt understand how to re-arrange the equation

myininaya (myininaya):

your 2nd problem had a bit more steps than 1st problem in the problem we had to undo alot of stuff to get to i

OpenStudy (anonymous):

thank you!

myininaya (myininaya):

first the i was being added to the 1 and then that was to the nth power but even all of that was being multiplied by that P so we would from outside to inside to get that i we had to undo the multiplication of P by dividing by P on both side we had to get rid of the nth power by raising both sides to the 1/n th power and then finally to get rid of the adding 1 we had to subtract 1 on both sides

myininaya (myininaya):

your welcome

OpenStudy (anonymous):

so basically to solve for P i use P = A / (1 + i)^n and to solve for i, i use i = (A/P)^(1/n) - 1

myininaya (myininaya):

well yeah you already solved for P and i i think you are asking me to confirm your answers and yes that is right those are your answers

OpenStudy (anonymous):

THANKS :)

OpenStudy (anonymous):

i understand it now :)